Date: Mon, 18 Jan 1999 16:28:06 -0800 (PST)
To: 2mass-staff
Subject: IPAC 2MASS WG Meeting #176 Minutes
Cc: chas,, bgreen,
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: vbvIbJk2g+Bh2TbO/ifsdQ==

           IPAC 2MASS Working Group Meeting #176 Minutes

Attendees: R. Cutri, S. Van Dyk, D. Kirkpatrick, R. Tam, H. McCallon, 
           D. Engler, B. Nelson, J. White, W. Wheaton, L. Fullmer, 
           R. Beck, G. Kopan [at end]


1.) D. Kirkpatrick Runs

2.) AAS Meeting in Austin

3.) 2MASS Sampler

4.) Spring Data Release

5.) Other Project Items

6.) New and Standing Action Items


1.) D. Kirkpatrick Runs

     D. Kirkpatrick returned recently from three very successful observing runs
on Mauna Kea: one run on Keck, another at UKIRT, and a third again on Keck.  All
nights were great, providing a perfect nine out of nine nights.  The number of
L dwarfs has now increased from 36 to 71.  He was primarily looking for those
stars that are cooler than cool, namely the methane T dwarfs.  He has not found
one Gl 229B counterpart in ~3000 square degrees.

2.) AAS Meeting in Austin

     S. Van Dyk is getting the exhibit produced for the AAS meeting.  G. Kopan
was not present at the meeting, since he was going to JPL to pick up the large
Galactic Center panel for the exhibit.  [G. Kopan later arrived near the very
end of the meeting; the panel was unfurled and inspected, before being packed
for shipping.]  S. Van Dyk had picked up the center panel at Graphics Vending
at JPL.  They had gotten it right on the first try, and it was packed up for 
shipping to Austin earlier in the day.  R. Cutri recommended to S. Van Dyk to 
make sure that Science Team members attending the meeting were contacted 
regarding staffing the exhibit during the meeting.

3.) 2MASS Sampler
     The Sampler is now online via the IRSA web interface.  An initial small
email exploder had been sent out, notifying about 250 astronomers of the 
availability of the data release.  The CD-ROM production is still tied up with
the vendor.  It is not clear yet exactly on which day the CD-ROMs will arrive
at the AAS meeting, to be distributed there at the exhibit.  2000 copies of the
CD-ROM have been ordered.

4.) Spring Data Release

     Telecons have been taking place between UMass and IPAC, discussing the
criteria for the Spring release.  Some of the issues include what will be the
minimum quality requirements for data to get into the catalogs.  What will be
the criteria for image elongation and maximum seeing, for instance?  Also being
discussed is the issue of untracked seeing, background residuals, and positional
quality.  One question is, do we release the "ACT-challenged" data?

     The Spring release will consist of about 4500 tiles, cutting the Galactic
Plane a few times.  A first cut of the catalogs must be made by mid-January.
All 2MASS'ers will be put to work.

     The issue of how to deal with redundant sources in scan overlap regions has
been discussed.  Do we merge the sources, using a weighted average, or use some
geometrical criterion, say, choosing either the left or right duplicate?  The
second option has been chosen.  Only one source apparition will be selected,
with the apparition closest to the scan center being chosen.

     The data release is in only 2 1/2 months.  It must be frozen in place in
1 1/2 months, to allow a month's worth of testing.  

     Future discussions in telecons will take place Tuesday mornings at 9 AM 

5.) Other Project Items

     W. Wheaton discussed PSF generation.  The first of two nights of testing
has been done.  The new PSFs will be for 980909s to the present.  The lookup
tables need to be ready for the pipeline.  Getting the next round of PSFs is
important, since the data with the new Leach electronics need to be scrutinized.
R. Cutri asked if we need more human resources.  W. Wheaton indicated probably
not; any delay in the PSF generation is probably related to the software.  The
routine "PSFMake" is currently rejecting more stars that it should be, so the
routine needs a little tweaking.  K. Marsh will be brought back on to help in
January.  W. Wheaton indicated that some scans in the recent nights have only
a few stars with which to work.

[CORRECTION:  The PSFs in question were for the interval 980605s to 980919s.
This time interval is currently being processed on barney, as this correction
to the notes is being written.  The results up to 980919s will be analyzed and
evaluated by W. Wheaton.]

     R. Cutri asked D. Kirkpatrick and B. Nelson about the QA backlog.  They
reported that we are currently 19 nights behind processing.  The QA'ers will be
helped a bit by the short-term shutdown of barney.  R. Cutri urged the QA
scientists to harrass UMass to get more timely approvals or rejections.

     G. Kopan discussed the K-only offset, which needs to be rectified.  The
question was raised as to whether the PSF selection is toggling properly for 
the seeing.  Pixphot seems to be doing better for the holes in H, but there are
still some depressions in K.
6.) New and Standing Action Items

%                                                      %
%                 ACTION ITEMS                         %
%                                                      %

Team members are urged to have a Happy New Year.

(Standing Action Items:)

a) Team members should continue to analyze data for Spring data release.  
Members will likely be called upon to assist in steps toward catalog generation
for the release.